# Aaronson-Ambainis Conjecture Is True For Random Restrictions

### 21 Feb 2024 v1

@@ 1. Introduction
One of the central open problems in the field of quantum query complexity is finding if there exists a partial function which is defined on a large fraction of the Boolean hypercube (say, constant) but whose quantum query complexity and classical query complexity are super-polynomially separated. The seminal result of Beals, Burhman et al. [4] shows that no such separation is possible when the function is defined on the entire hypercube. On the other hand, functions for which we know such a separation (e.g. - Forrelation [1] , Bernstein-Vazirani [5]) are defined on an exponentially small fraction of the hypercube. A possible explanation as to why all known functions exhibiting large gaps between quantum and classical query complexity have very small support size would be the following folklore conjecture:
@@
Aaronson and Ambainis [2] proposed the following query algorithm to estimate $P$: suppose the variance of the function is sufficiently small. Then we terminate the query algorithm and output the average over the unqueried coordinates. If not, we query the coordinate with the highest influence and restrict the function according to the response received. We keep doing this until we have made too many queries or the variance has become sufficiently low. In order to show that this algorithm gives an accurate estimate, [2] observed that it is sufficient to prove the following conjecture.
@@
As a side remark, we mention that O’Donnell et al. [16] had shown previously that functions which can be approximated by decision trees have a coordinate with high influence. So conjectures 1.2 and 1.3 are equivalent.
Aaronson-Ambainis conjecture has received significant attention in the past few years. A 2006 result of Dinur, Friedgut et al. [8] shows that the conjecture is true if $\mathsf{poly}(d)$ is replaced by $\exp(d)$. In 2012, Montanaro [14] proved the conjecture in the special case of block-multilinear forms where all coefficients have the same magnitude. In 2016, O’Donnell and Zhao [17] showed that it suffices to prove the conjecture for a special class of polynomials known as one-block decoupled polynomials. In 2020, Keller and Klein [11] claimed to have found a proof for the conjecture but their paper had a subtle flaw and turned out to be wrong. More recently, Lovett and Zhang [13] initiated a new line of attack using the notions of fractional block sensitivity and fractional certificate complexity. In 2022, Bansal, Sinha, Wolf [3] proved that this conjecture is true for completely bounded block multilinear forms - a class of polynomials that captures a special kind of quantum query algorithms.
@@ 3. Notations and preliminaries
    A classical query algorithm $A$ (or equivalently, a decision tree) for computing a function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ can access the input $x\in\{\pm 1\}^{n}$ by adaptively issuing queries to its bits. We assume internal computations have no cost. The depth of the query algorithm/decision tree is the maximum number of bit queries issued on an input. We say $A$ $\epsilon$-approximates $f$ if $||f-A||_{2}^{2}=\mathsf{E}_{x\in\{\pm 1\}^{n}}\left[(f(x)-A(x))^{2}\right]\leq\epsilon$.
@@
    A quantum query algorithm can access the input $x\in\{\pm 1\}^{n}$ via an oracle $O_{x}$. The register has $n$ qubits and some ancilla qubits. The oracle $O_{x}$ applies the following unitary operation on the first $n$ qubits:
    $O_{x}\left\lvert s_{1}s_{2}\cdots s_{n}\right\rangle=(-1)^{\langle x,s\rangle}\left\lvert s_{1}s_{2}\cdots s_{n}\right\rangle$
@@
    $\Pr[Q\text{ outputs }1\text{ on input }x]\begin{cases}\geq 2/3&\text{ if }f(x)=1\\ \leq 1/3&\text{ if }f(x)=0\end{cases}$
@@
In this section we recall some results from analysis of boolean functions. A good reference is O’ Donnell’s textbook [15].
@@
    Given two functions $f,g\{\pm 1\}^{n}\rightarrow\mathbb{R}$, we say $g$ $\epsilon-$approximates $f$ if $||f-g||_{2}^{2}=\mathsf{E}_{x}\left[(f(x)-g(x))^{2}\right]\leq\epsilon$.
@@
    For a point $x\in\{\pm 1\}^{n}$ and a subset $S\subseteq[n]$ and $-1\leq\rho\leq 1$, we define a distribution $N_{\rho,S}(x)$ on {±$\{\pm 1\}^{n}$ as follows:
@@
    For $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ and $-1\leq\rho\leq 1$ define $T_{\rho}f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ by
@@
    A function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ is a junta of arity $l$ or $l$-junta if there exists a subset $S\subseteq[n],|S|\leq l$ such that $f$ only depends on the coordinates in $S$. We say $f$ is a $(\epsilon,l)$ junta if it can be $\epsilon$-approximated by a $l$-junta, i.e., there exists a $l$-junta $g$ such that $||f-g||_{2}^{2}\leq\epsilon$.
@@
[…]
@@ 4. Proof Overview
The main technical tool in this paper is a structural result for bounded low-degree functions similar in spirit to Hastad’s switching lemma [10]. Let $f:\{\pm 1\}^{n}\rightarrow[0,1]$ be a polynomial of degree $d$, and let $\rho$ denote a random restriction with survival probability $\dfrac{\log(d)}{Cd}$. We show that for some constant $C$,
@@
Now we give a brief overview of how we prove the switching lemma. The starting point is the work by Dinur, Friedgut et al. [8] which states the following:
@@
In order to make this approach work, we need to improve the tail bound $\exp(-O(k^{2}\log(k))/\epsilon)$. The problematic term is the quadratic $\exp(-O(k^{2}))$ in the exponential. If the dominant term were to the order of $\exp(-O(k))$ instead, the calculations would go through. Can we hope to increase the tail bound to $\exp(-O(k))$ while paying a cost by increasing the junta arity? Unfortunately, again, this is not possible: [8] constructs a function which shows that the tail bound is essentially tight upto the $\log(k)$ factor - their function has $||f^{>k}||_{2}^{2}\approx\exp(-\Theta(k^{2}))$ but approximating it to even $1/3$ accuracy requires reading $\Omega(n)$ coordinates. Our key observation is that the function constructed by [8] has full degree whereas we are working with random restrictions of a low degree function, so in addition to the fact that Fourier tail of $f_{\rho}$ above level $\log(d)$ is very small, we also know that $f_{\rho}$ has degree $d$. Can we hope to improve the tail bound in Theorem 4.1 if we have the additional restriction that the function is of degree $d$? Indeed, this turns out to be true. We prove the following result in Theorem 4.2:.
@@
Below we briefly discuss how we are able to improve the tail bound under the additional degree assumption. Dinur et al [8] prove their tail bound by showing the following result (we are omitting the exact quantitative parameters here for reading convenience).
@@
To prove Theorem 4.1, [8] applies Theorem 4.3 on the truncated function $h=f^{\leq k}=\displaystyle\sum_{|S|\leq k}\hat{f}(S)\chi_{S}$ and takes $g$ to be the original function $f$. This then lower bounds $E[(f-f^{\leq k})^{2}]$ which is precisely the Fourier tail above weight $k$. Thus, the distance lower bound $\varepsilon$ in Theorem 4.3 governs the Fourier tail lower bound in Theorem 4.1. Since we have the additional information that $f$ is of degree $d$, for our purposes it will suffice to bound the distance from bounded degree $d$ functions, not necessarily all bounded functions. In Theorem 6.1 we prove a result of the following form (again, we are omitting the exact parameters for reading convenience)
@@
In order to prove Theorem 4.4 we shall use the main idea of the proof of [8] along with a structural restriction for bounded low-degree functions discovered first in [4]. Given any function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ and $x\in\{\pm 1\}^{n}$ define the block sensitivity of $f$ at $x$ to be
@@
[…]
We give a high-level overview of how we are able to improve upon the bound of $\varepsilon$ using the fact that the block sensitivity of a bounded low degree function is small. At one point in their proof, [8] lower bounds the probability of a linear form of Rademacher random variables $l(x_{1},x_{2},\cdots,x_{t})=a_{1}x_{1}+\cdots+a_{t}x_{t}$ exceeding a certain threshold times its standard deviation, i.e.,
@@
For each such point $x$ where this linear form is high, [8] shows that many related points $x^{\prime}$ must have $f(x^{\prime})>2$. Using this they conclude that $f$ must deviate from the interval $[0,1]$ too often and therefore cannot be approximated by any bounded function.
We follow the proof of [8] up until this point. Instead of directly lower bounding the probability that $a_{1}x_{1}+\cdots+a_{t}x_{t}$ is high, we partition the set of variables $[t]$ into $L$ blocks $B_{1},\cdots,B_{L}$ ($L$ is an appropriately chosen parameter) such that each block gets roughly same total weight: for all $j\in[L],$
@@
Our advantage is that we need to set $\alpha^{\prime}$ so that we can conclude $|f(z)-f(z^{\prime})|$ is only somewhat larger than $\Omega(d^{2}/L)$ (as opposed to $\Omega(1)$ in [8]) - by setting $L$ large enough this allows us to set a much smaller $\alpha$ and get rid of the quadratic exponential dependence.
@@ 5. Tools
The proof of these lemmas can be found in [8].
@@
This is the main result of [8]. We use a slight variant. First we recall some known results from approximation theory.
@@
For any $k$, there exist constants $\rho_{1},\rho_{2},\cdots,\rho_{k+1}\in[-1/2,1/2]$ with the following property: for any polynomial of degree $k$, $p(x)=a_{0}+a_{1}x+\cdots+a_{k}x^{k}$, there exists a $j\in[k+1]$ such that $p(\rho_{j})\geq\dfrac{a_{1}}{2(k+1)}.$
@@
$\Pr\left[f(z)-f(x_{0})\geq\dfrac{\gamma}{2(k+1)}\right]\geq\dfrac{1}{(k+1)B^{k}}.$
@@
$\Pr_{z\leftarrow N_{\rho}(x_{0})}\left[f(z)-f(x_{0})\geq\dfrac{\gamma}{2(k+1)}\big{|}\rho=\rho_{h}\right]\geq\dfrac{1}{B^{k}}.$
@@
$\Pr_{z\leftarrow N_{\rho}(x_{0})}\left[f(z)-f(x_{0})\geq\dfrac{\gamma}{2(k+1)}\right]\geq\dfrac{1}{(k+1)B^{k}}.$
@@
reduces by $\min\{a_{i}|a_{i}\neq 0\}$ at each step, so at some point of time it must be 0 at which point the algorithm terminates and returns a valid partition.
@@ 6.1. Improved tail bound for low degree functions
For a subset $J\subseteq[n]$, consider the junta $u:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ which reads the coordinates of $J$ and outputs the average over the unqueried coordinates. It is easy to see that $u(x)=\displaystyle\sum_{S\subseteq J}\hat{f}(S)\chi_{S}(x),$ so $||u-f||_{2}^{2}=\displaystyle\sum_{S\not\subseteq J}\hat{f}(S)^{2}.$ Thus, $u$ approximates $f$ if and only if $\sum_{S\not\subseteq J}\hat{f}(S)^{2}$ is small.
@@
In fact, it is easy to see that there exists a junta $u$ depending only on coordinates of $J$ such that $||f-u||_{2}^{2}\leq\epsilon$ if and only if $\sum_{S\not\subseteq J}\hat{f}(S)^{2}\leq\epsilon$. This immediately follows from the Fourier expansion of $f-u$.
@@
Suppose, for the sake of contradiction, there exists a degree $d$ polynomial $g:\{\pm 1\}^{n}\rightarrow[0,1]$ such that $\mathsf{E}[(f(x)-g(x))^{2}]\leq\delta$. Throughout the rest of the proof, for a string $s_{1}\in\{\pm 1\}^{[n]\setminus U}$ and a string $s_{2}\in\{\pm 1\}^{U}$, the pair $(s_{1},s_{2})$ denotes the string $s\in\{\pm 1\}^{n}$ which agrees with $s_{1}$ on $[n]\setminus U$ and with $s_{2}$ on $U$. Consider the following randomized procedure which returns a real number.
Procedure 1: (1) Sample a $y\in\mathsf{GOOD}$ uniformly at random. (2) Sample $\rho\leftarrow\{\rho_{1},\rho_{2},\cdots,\rho_{k+1}\}$ uniformly at random. (3) Sample $z\leftarrow\{\pm 1\}^{U}$ uniformly at random. (4) Let $\mathsf{DIVIDE}(y)=(B_{1},B_{2},\cdots,B_{L})$. Sample $\tilde{z}^{(i)}\leftarrow N_{B_{i},\rho}(z)$ for $1\leq i\leq L$. (5) Return $\sum_{i=1}^{L}|f(y,z)-f(y,\tilde{z}^{(i)})|$. We estimate the probability that procedure 1 returns a number $>15d^{2}$ in two different ways. First, we obtain a lower bound from our assumption that $\sum_{S\not\subseteq J}\hat{f}(S)^{2}\geq\mu$. Then, we obtain an upper bound from the assumption that $\mathsf{E}[(f(x)-g(x))^{2}]\leq\delta$. These two bounds will contradict each other - and that will prove the theorem.
@@
Now consider a uniformly sampled $y\in\mathsf{GOOD}$. Observe that as we sample $z\leftarrow\{\pm 1\}^{U}$ u.a.r, $\rho\leftarrow\{\rho_{1},\cdots,\rho_{k+1}\}$ u.a.r and $\tilde{z}^{(i)}\leftarrow N_{\rho,B_{i}}(z)$, the marginal distribution of $\tilde{z}^{(i)}$ is uniform on {±$\{\pm 1\}^{U}$. By Markov’s inequality, we have
@@
By union bound, the probability that $(f(y,z)-g(y,z))^{2}\geq\dfrac{1}{L^{2}}$ or for some $i$, $(f(y,\tilde{z}^{(i)})-g(y,\tilde{z}^{(i)}))^{2}\geq\dfrac{1}{L^{2}}$ is at most $(L+1)\dfrac{L^{2}\delta}{\mu}(3W)^{k}\leq\dfrac{2L^{3}\delta}{\mu}(3W)^{k}$. Our choice of $\delta$ ensures that this quantity is less than $<\dfrac{1}{2L\times(2B)^{k}}$. Observe that if none of these bad events holds, since the block sensitivity of $g$ is bounded above by $6d^{2}$ (Theorem 4.5), we have that
@@
As promised, we get conflicting lower and upper bounds for the probability that procedure 1 returns a number $>15d^{2}$. This is our desired contradiction. ∎
Now we show that we can improve the tail bound of [8] under the additional assumption that $f$ has low degree. This follows straightforwardly from Theorem 6.1.

### 04 Mar 2026 v2

@@ 1. Introduction
One of the central open problems in the field of quantum query complexity is finding if there exists a partial function which is defined on a large fraction of the Boolean hypercube (say, constant) but whose quantum query complexity and classical query complexity are super-polynomially separated. The seminal result of Beals, Burhman et al. [BBC+98] shows that no such separation is possible when the function is defined on the entire hypercube. On the other hand, functions for which we know such a separation (e.g. - Forrelation [AA18] , Bernstein-Vazirani [BV97]) are defined on an exponentially small fraction of the hypercube. A possible explanation as to why all known functions exhibiting large gaps between quantum and classical query complexity have very small support size would be the following folklore conjecture:
@@
Aaronson and Ambainis [AA11] proposed the following query algorithm to estimate $P$: suppose the variance of the function is sufficiently small. Then we terminate the query algorithm and output the average over the unqueried coordinates. If not, we query the coordinate with the highest influence and restrict the function according to the response received. We keep doing this until we have made too many queries or the variance has become sufficiently low. In order to show that this algorithm gives an accurate estimate, [AA11] observed that it is sufficient to prove the following conjecture.
@@
As a side remark, we mention that O’Donnell et al. [OSSS05] had shown previously that functions which can be approximated by decision trees have a coordinate with high influence. So conjectures 1.2 and 1.3 are equivalent.
Aaronson-Ambainis conjecture has received significant attention in the past few years. A 2006 result of Dinur, Friedgut, Kindler, O’Donnell [DFKO06] shows that the conjecture is true if $\mathsf{poly}(d)$ is replaced by $\exp(d)$. In 2012, Montanaro [Mon12] proved the conjecture in the special case of block-multilinear forms where all coefficients have the same magnitude. In 2016, O’Donnell and Zhao [OZ16] showed that it suffices to prove the conjecture for a special class of polynomials known as one-block decoupled polynomials. In 2020, Keller and Klein [NK19] claimed to have found a proof for the conjecture but their paper had a subtle flaw and turned out to be wrong. More recently, Lovett and Zhang [LZ23] initiated a new line of attack using the notions of fractional block sensitivity and fractional certificate complexity. In 2022, Bansal, Sinha, Wolf [BSdW22] proved that this conjecture is true for completely bounded block multilinear forms - a class of polynomials that captures a special kind of quantum query algorithms.
@@ 3. Notations and preliminaries
    A classical query algorithm $A$ (or equivalently, a decision tree) for computing a function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ can access the input $x\in\{\pm 1\}^{n}$ by adaptively issuing queries to its bits. We assume internal computations have no cost. The depth of the query algorithm/decision tree is the maximum number of bit queries issued on an input. We say AA $\epsilon$-approximates $f$ if $f-A||_{2}^{2}=\mathsf{E}_{x\in\{\pm 1\}^{n}}\left[(f(x)-A(x))^{2}\right]\leq\epsilon$.
@@
    A quantum query algorithm can access the input $x\in\{\pm 1\}^{n}$ via an oracle $O_{x}$. The oracle acts on a fixed set of $\lceil\log(n)\rceil$ qubits (which the query algorithm has access to) in the following manner:
    $O_{x}\lvert j\rangle=(-1)^{x_{j}}\lvert j\rangle\text{ for all }j\in[n].$
@@
[…]
@@
In this section we recall some results from analysis of boolean functions. A good reference is O’ Donnell’s textbook [O’D14].
@@
    Given two functions $f,g\{\pm 1\}^{n}\rightarrow\mathbb{R}$, we say $g$ $\epsilon$-approximates $f$ if $f-g||_{2}^{2}=\mathsf{E}_{x}\left[(f(x)-g(x))^{2}\right]\leq\epsilon$.
@@
    For a point $x\in\{\pm 1\}^{n}$ and a subset $S\subseteq[n]$ and $1\leq\rho\leq 1$, we define a distribution $N_{\rho,S}(x)$ on $\{\pm 1\}^{n}$ as follows:
@@
    For $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ and $1\leq\rho\leq 1$ define $T_{\rho}f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ by
@@
    A function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ is a junta of arity $l$ or $l$-junta if there exists a subset $S\subseteq[n],|S|\leq l$ such that $f$ only depends on the coordinates in $S$. We say $f$ is a $(\epsilon,l)$ junta if it can be $\epsilon$-approximated by a $l$-junta, i.e., there exists a $l$-junta $g$ such that $f-g||_{2}^{2}\leq\epsilon$.
@@
Throughout the paper, all growing parameters (e.g., the degree $d$) will be assumed to be larger than some sufficiently big constant. This is to make the expressions look neat, as we will be replacing terms like $(C_{1})^{k}\text{poly}(k)$ by $(C_{2})^{k}$ where $C_{2}>C_{1}$.
@@ 4. Proof Overview
The main technical tool in this paper is a structural result for bounded low-degree functions similar in spirit to Hastad’s switching lemma [Hås86]. Let $f:\{\pm 1\}^{n}\rightarrow[0,1]$ be a polynomial of degree $d$, and let $\rho$ denote a random restriction with survival probability $\dfrac{\log(d)}{Cd}$. We show that for some constant $C$,
@@
Now we give a brief overview of how we prove the switching lemma. The starting point is the work by Dinur, Friedgut et al. [DFKO06] which states the following:
@@
In order to make this approach work, we need to improve the tail bound $\exp(-O(k^{2}\log(k))/\epsilon)$. The problematic term is the quadratic $\exp(-O(k^{2}))$ in the exponential. If the dominant term were to the order of $\exp(-O(k))$ instead, the calculations would go through. Can we hope to increase the tail bound to $\exp(-O(k))$ while paying a cost by increasing the junta arity? Unfortunately, again, this is not possible: [DFKO06] constructs a function which shows that the tail bound is essentially tight upto the $\log(k)$ factor - their function has $f^{>k}||_{2}^{2}\approx\exp(-\Theta(k^{2}))$ but approximating it to even $1/3$ accuracy requires reading $\Omega(n)$ coordinates. Our key observation is that the function constructed by [DFKO06] has full degree whereas we are working with random restrictions of a low degree function, so in addition to the fact that Fourier tail of $f_{\rho}$ above level $\log(d)$ is very small, we also know that $f_{\rho}$ has degree $d$. Can we hope to improve the tail bound in Theorem 4.1 if we have the additional restriction that the function is of degree $d$? Indeed, this turns out to be true. We prove the following result in Theorem 4.2:.
@@
Below we briefly discuss how we are able to improve the tail bound under the additional degree assumption. Dinur, Friedgut, Kindler, O’Donnell [DFKO06] prove their tail bound by showing the following result (we are omitting the exact quantitative parameters here for reading convenience).
@@
To prove Theorem 4.1, [DFKO06] applies Theorem 4.3 on the truncated function $h=f^{\leq k}=\displaystyle\sum_{|S|\leq k}\hat{f}(S)\chi_{S}$ and takes $g$ to be the original function $f$. This then lower bounds $E[(f-f^{\leq k})^{2}]$ which is precisely the Fourier tail above weight $k$. Thus, the distance lower bound $\varepsilon$ in Theorem 4.3 governs the Fourier tail lower bound in Theorem 4.1. Since we have the additional information that $f$ is of degree $d$, for our purposes it will suffice to bound the distance from bounded degree $d$ functions, not necessarily all bounded functions. In Theorem 6.1 we prove a result of the following form (again, we are omitting the exact parameters for reading convenience)
@@
In order to prove Theorem 4.4 we shall use the main idea of the proof of [DFKO06] along with a structural restriction for bounded low-degree functions discovered first in [BBC+98]. Given any function $f:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ and $x\in\{\pm 1\}^{n}$ define the block sensitivity of $f$ at $x$ to be
@@
[BBC+98] If $f:\{\pm 1\}^{n}\rightarrow[0,1]$ has degree $d$, $\text{bs}(f)\leq 6d^{2}$.
We give a high-level overview of how we are able to improve upon the bound of $\varepsilon$ using the fact that the block sensitivity of a bounded low degree function is small. At one point in their proof, [DFKO06] lower bounds the probability of a linear form of Rademacher random variables $l(x_{1},x_{2},\cdots,x_{t})=a_{1}x_{1}+\cdots+a_{t}x_{t}$ exceeding a certain threshold times its standard deviation, i.e.,
@@
For each such point $x$ where this linear form is high, [DFKO06] shows that many related points $x^{\prime}$ must have $f(x^{\prime})>2$. Using this they conclude that $f$ must deviate from the interval $[0,1]$ too often and therefore cannot be approximated by any bounded function.
We follow the proof of [DFKO06] up until this point. Instead of directly lower bounding the probability that $a_{1}x_{1}+\cdots+a_{t}x_{t}$ is high, we partition the set of variables $[t]$ into $L$ blocks $B_{1},\cdots,B_{L}$ ($L$ is an appropriately chosen parameter) such that each block gets roughly same total weight: for all $j\in[L],$
@@
Our advantage is that we need to set $\alpha^{\prime}$ so that we can conclude $|f(z)-f(z^{\prime})|$ is only somewhat larger than $\Omega(d^{2}/L)$ (as opposed to $\Omega(1)$ in [DFKO06]) - by setting $L$ large enough this allows us to set a much smaller $\alpha$ and get rid of the quadratic exponential dependence.
@@ 5. Tools
The proof of these lemmas can be found in [DFKO06].
@@
This is the main result of [DFKO06]. We use a slight variant. First we recall some known results from approximation theory.
@@
For any $k$, there exist constants $\rho_{1},\rho_{2},\cdots,\rho_{k+1}\in[-1/2,1/2]$ with the following property: for any polynomial of degree $k$, $p(x)=a_{0}+a_{1}x+\cdots+a_{k}x^{k}$, there exists a $j\in[k+1]$ such that $|p(\rho_{j})|\geq\dfrac{|a_{1}|}{2(k+1)}.$
@@
$\Pr\left[\left|f(z)-f(x_{0})\right|\geq\dfrac{\gamma}{2(k+1)}\right]\geq\dfrac{1}{(k+1)B^{k}}.$
@@
$\Pr_{z\leftarrow N_{\rho}(x_{0})}\left[\left|f(z)-f(x_{0})\right|\geq\dfrac{\gamma}{2(k+1)}\big|\rho=\rho_{h}\right]\geq\dfrac{1}{B^{k}}.$
@@
$\Pr_{z\leftarrow N_{\rho}(x_{0})}\left[\left|f(z)-f(x_{0})\right|\geq\dfrac{\gamma}{2(k+1)}\right]\geq\dfrac{1}{(k+1)B^{k}}.$
@@
reduces by $\min\{a_{i}|a_{i}\neq 0\}/2L$ at each step, so at some point of time it must be 0 at which point the algorithm terminates and returns a valid partition.
@@ 6.1. Improved tail bound for low degree functions
For a subset $J\subseteq[n]$, consider the junta $u:\{\pm 1\}^{n}\rightarrow\mathbb{R}$ which reads the coordinates of $J$ and outputs the average over the unqueried coordinates. It is easy to see that $u(x)=\displaystyle\sum_{S\subseteq J}\hat{f}(S)\chi_{S}(x),$ so $u-f||_{2}^{2}=\displaystyle\sum_{S\not\subseteq J}\hat{f}(S)^{2}.$ Thus, $u$ approximates $f$ if and only if $\sum_{S\not\subseteq J}\hat{f}(S)^{2}$ is small.
@@
In fact, it is easy to see that there exists a junta $u$ depending only on coordinates of $J$ such that $f-u||_{2}^{2}\leq\epsilon$ if and only if $\sum_{S\not\subseteq J}\hat{f}(S)^{2}\leq\epsilon$. This immediately follows from the Fourier expansion of fuf-u.
@@
Suppose, for the sake of contradiction, there exists a degree $d$ polynomial $g:\{\pm 1\}^{n}\rightarrow[0,1]$ such that $\mathsf{E}[(f(x)-g(x))^{2}]\leq\delta$. Throughout the rest of the proof, for a string $s_{1}\in\{\pm 1\}^{[n]\setminus U}$ and a string $s_{2}\in\{\pm 1\}^{U}$, the pair $(s_{1},s_{2})$ denotes the string $s\in\{\pm 1\}^{n}$ which agrees with $s_{1}$ on $n]\setminus U$ and with $s_{2}$ on $U$. Consider the following randomized procedure which returns a real number.
Procedure 1: (1) Sample a $y\in\mathsf{GOOD}$ uniformly at random. (2) Sample $\rho\leftarrow\{\rho_{1},\rho_{2},\cdots,\rho_{k+1}\}$ uniformly at random. (3) Sample $z\leftarrow\{\pm 1\}^{U}$ uniformly at random. (4) Let $\mathsf{DIVIDE}(y)=(B_{1},B_{2},\cdots,B_{L})$. Sample $\tilde{z}^{(i)}\leftarrow N_{B_{i},\rho}(z)$ for $1\leq i\leq L$. (5) Return $\sum_{i=1}^{L}|f(y,z)-f(y,\tilde{z}^{(i)})|$. We estimate the probability that procedure 1 returns a number $15d^{2}$ in two different ways. First, we obtain a lower bound from the definition of GOOD. Then, we obtain an upper bound from the assumption that $\mathsf{E}[(f(x)-g(x))^{2}]\leq\delta$ and the fact that we have a lower bound on $\text{Pr}_{y}[y\in\text{GOOD}]$ (which, recall, follows from the assumption that $\sum_{S\not\subseteq J}\hat{f}(S)^{2}\geq\mu$). These two bounds will contradict each other - and that will prove the theorem.
@@
Now consider a uniformly sampled $y\in\mathsf{GOOD}$. Observe that as we sample $z\leftarrow\{\pm 1\}^{U}$ u.a.r, $\rho\leftarrow\{\rho_{1},\cdots,\rho_{k+1}\}$ u.a.r and $\tilde{z}^{(i)}\leftarrow N_{\rho,B_{i}}(z)$, the marginal distribution of $\tilde{z}^{(i)}$ is uniform on $\{\pm 1\}^{U}$. By Markov’s inequality, we have
@@
By union bound, the probability that $(f(y,z)-g(y,z))^{2}\geq\dfrac{1}{L^{2}}$ or for some $i$, $(f(y,\tilde{z}^{(i)})-g(y,\tilde{z}^{(i)}))^{2}\geq\dfrac{1}{L^{2}}$ is at most $(L+1)\dfrac{L^{2}\delta}{\mu}(3W)^{k}\leq\dfrac{2L^{3}\delta}{\mu}(3W)^{k}$. Our choice of $\delta$ ensures that this quantity is less than $\dfrac{1}{2L\times(2B)^{k}}$. Observe that if none of these bad events holds, since the block sensitivity of $g$ is bounded above by $6d^{2}$ (Theorem 4.5), we have that
@@
As promised, we get conflicting lower and upper bounds for the probability that procedure 1 returns a number $15d^{2}$. This is our desired contradiction. ∎
Now we show that we can improve the tail bound of [DFKO06] under the additional assumption that $f$ has low degree. This follows straightforwardly from Theorem 6.1.